OpenStudy (andu1854):

lim 1/b - 1/2 b->2 -------- b-2 Find indicated limit

ganeshie8 (ganeshie8):

combine the fractions in numerator and you will see something cancels out

OpenStudy (anonymous):

i know that this is not an infinite limit

OpenStudy (anonymous):

i think if i did it correctly you get 0/0 making it not an infinite limit

ganeshie8 (ganeshie8):

@andu1854 still here ? :)

OpenStudy (anonymous):

\[\frac{\frac{1}{b}-\frac{1}{2}}{b-2}=-\frac{1}{2 b} \]

OpenStudy (anonymous):

Limit is\[-\frac{1}{4} \]

OpenStudy (andu1854):

Thank you very much guys...